# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ 4 LATS AND LAGEX: ALGORITHM DESIGN AND ANALYSIS
11:     Update loss for regret minimizer: Update with $L_{t}=\langle\boldsymbol{\omega}_{t}^{\star},U_{t}\rangle-{\boldsymbol{l}_{t}^{\star}}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}^{\star}$
12:     t = t + 1
13:end while
14:Recommended policy:$\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\hat{{\boldsymbol{\mu}}}_{t}^{\top}\boldsymbol{\pi}$
Component 2: Recommendation rule. Once the stopping rule is fired, the agent recommends a policy based on the current estimate of $\hat{{\boldsymbol{\mu}}}_{t}$ according the rule $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}_{\tau_{\delta}}}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{\tau_{\delta}}}^{r}}\hat{{\boldsymbol{\mu}}}_{t}^{\top}\boldsymbol{\pi}$.
a. LATS. The algorithm LATS (Algorithm 1) uses a Track and Stop strategy adapted to the unknown constraint setting. We use blue markers in the pseudocode (Algorithm 1) to denote the novel approaches taken to handle the challenge of estimating the feasible space per step. The algorithm first warms up the parameter estimates by playing each arm once. Then until the test statistic jumps over the threshold, first it computes the optimal allocation (Line 4) under the estimated feasible space at the current step solving the Lagrangian relaxed optimization problem in Proposition 1 by plugging in the best choice of Lagrangian multiplier optimized in the previous iteration. Using this current optimal allocation we optimize the Lagrangian multiplier maintaining the bounds of it’s 1-norm stated in Theorem 1. The algorithm uses C-tracking (Garivier and Kaufmann,, 2016) to track the action taken per step. Finally we observe the instantaneous reward and cost feedback to update the parameter estimates based on them (Line 9).
Theorem 6.
Let $\mathfrak{s}$ be the shadow price $\mathfrak{s}\mathrel{\triangleq}\frac{\Gamma_{\mathrm{max}}}{\Gamma_{\mathrm{min}}}$ of the slack $\Gamma$. For any $\alpha>1$, the expected stopping time of LATS satisfies $\lim_{\begin{subarray}{c}\delta\rightarrow 0\end{subarray}}\frac{\mathbb{E}[\tau_{\delta}]}{\log(1/\delta)}\leq\alpha T_{\mathcal{F},r}({\boldsymbol{\mu}})(1+\mathfrak{s}).$ (Proof is in Appendix G.3.)
@@
b. LAGEX. Algorithms based on track and stop mechanism tend to fail in case of larger problems where efficient optimization becomes a challenge due to the use of a max-min oracle per step. To improve on this we leverage the two-player zero sum game approach introduced in Degenne et al., 2019b . Algorithm 2 also starts by playing each arm once to warm up parameter estimates. Then it uses a allocation player (We have used AdaGrad) to optimize the allocation $\boldsymbol{\omega}_{t}$ in Line 5 against the most confusing instance w.r.t current estimate of ${\boldsymbol{\mu}}$ optimised by a instance player which minimizes $\sum_{a=1}^{K}\boldsymbol{\omega}_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-{\boldsymbol{l}_{t}^{\star}}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$ with respect to ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}({\boldsymbol{\mu}},\boldsymbol{\pi})$. Since our search space in closed and convex, the allocation player enjoys sub-linear regret of order $\mathcal{O}(\sqrt{t\log t})$, whereas the instance player computes the best confusing instance using the Lagrangian formulation of the weighted projection lemma stated in Proposition 1. Then in Line 11, Adagrad loss function is updated with a loss by introducing optimism as $U_{t}$ defined in Line 10. LAGEX also uses C-tracking as LATS to track the actions taken per step. Finally, LAGEX observes the instantaneous reward and cost to update the estimates.
Theorem 7.
The expected sample complexity of LAGEX satisfies
$\mathbb{E}[\tau]\leq T_{0}(\delta)+2\mathfrak{s}+C{K}\,.$
where $T_{0}(\delta)\mathrel{\triangleq}\max\left\{t\in\mathbb{N}:t\leq T_{\mathcal{F}}({\boldsymbol{\mu}})\left(\beta(t,\delta)+\mathcal{O}\left(\sqrt{t\log t}\right)\right)\right\}$ and $C\approx 21$ is a constant.
Implications. Theorem 6 and 7 show both LATS and LAGEX are asymptotically stable. Also, for LATS, the effect of the unknown constraints in the form of shadow price arises in a multiplicative way, whereas for LAGEX, it is additive. Thus, LAGEX should show a lower sample complexity than LATS, which is later validated in the experiments. The complete proof is in Appendix G.4.
@@ 5 EXPERIMENTAL ANALYSIS
Now, we empirically test performance of proposed algorithms and the baselines. We refer to Appendix I for the details about different setups, baseline algorithms, thresholds, and parameter values.
Figure 2: Lower bounds with and without constraints, and LAGEX for ${\boldsymbol{\mu}}_{3}\in[0.5,2.5]$ in Setup 1.
Figure 3: Sample complexity (median±$\pm$±std.) of algorithms for hard env. in Setup 1.
Figure 4: Sample complexity (median±$\pm$±std.) of algorithms for easy env. in Setup 1.
Figure 5: Sample complexity (median±$\pm$±std.) of algorithms for IMDB-50K.
Synthetic Data: Setup 1. We evaluate with two environments having means $[1.5,1.0,{\boldsymbol{\mu}}_{3},0.4,0.3,0.2,0.1]$.
Observation 1: Universality. We vary ${\boldsymbol{\mu}}_{3}$ from $0.5$ to $2.5$. For each environment, we plot the corresponding unconstrained BAI lower bounds (in red) and lower bounds under constraints (in blue) in Figure 5. We observe that the constraint problem gets easier with increasing ${\boldsymbol{\mu}}_{3}$. In contrast, the BAI problem changes non-monotonically. BAI problem gets harder when ${\boldsymbol{\mu}}_{3}$ is around $1.5$ as the suboptimality gap gets very small. But the constraint problem stays easier than BAI. In Figure 5, we also plot the median sample complexity of LAGEX across these environments over 500 runs. We observe that LAGEX grows parallel to the lower bound under constraints and can track it across environments.
Observation 2: Efficiency. We run all these algorithms in two environment: (i) hard with ${\boldsymbol{\mu}}_{3}=0.5$ and (ii) easy with ${\boldsymbol{\mu}}_{3}=1.3$. We call the first environment hard as it is harder than BAI and similarly, the second environment easy. In Figure 5 and 5 we observe that (i) among the algorithms with unknown constraints LAGEX incur the least sample complexity, and (ii) we pay a minimal cost than the known constraint Lagrangian algorithms in hard env whereas the price of estimating constraints is prominent in easy env.
Real Data: IMDB-50K Dataset. We evaluate LATS, LAGEX, and the baselines on IMDB 50K dataset (Maas et al.,, 2011). For ease of comparison, we use the same bandit environment as in (Carlsson et al.,, 2024). We use 12 movies. We search for the optimal policy which allocates weight at most $0.3$ to action movies and at least $0.3$ to family and drama movies. The true optimal policy is $[0.3,0.3,0,0,0.4,0,0,0,0,0,0,0]$. We assume $\delta=0.1$. We compare the same set of algorithms as Setup 1.
@@
6 DISCUSSION & FUTURE WORKS
We study the problem of pure exploration under unknown linear constraints. This problem requires tracking both mean vector and constraints to recommend a $r$-good feasible policy. We encompass this effect with a Lagrangian relaxation of the lower bound for known constraints. We further design an pessimistic estimate of the feasible set to ensure identification of the optimal feasible policy. These tools allows us to propose two algorithms LATS and LAGEX. We prove their sample complexity upper bounds, and conduct numerical experiments to find that LAGEX is the most efficient among baselines.
From the lower bound perspective, we have proposed a Lagrangian relaxation of the known lower bounds with plug-in estimates of the constraints. One might be interested to derive the lower bounds when both constraints and rewards are considered as feedback. In this paper, we focus on linear constraints on policies. One might be interested in non-linear constraints over the policy space. It would be interesting to extend our Lagrangian-based technique to nonlinear constraints.
Acknowledgments and Disclosure of Funding
We acknowledge the ANR JCJC project REPUBLIC (ANR-22-CE23-0003-01), the PEPR project FOUNDRY (ANR23-PEIA-0003), and the Inria-Japan associate team RELIANT for supporting the project.
@@ Appendix A Notations
${K}$ Number of Arms
$\mathbf{A}$ True constraint set
@@
$\mathcal{F}$ True feasible set w.r.t $\mathbf{A}$, =$\mathcal{F}=\{\mathbf{A}\in\mathbb{R}^{N\times K}:\mathbf{A}\boldsymbol{\pi}\leq 0\}$
$\tilde{\mathbf{A}}_{t}$ Pessimistic estimate of constraint set at time $t$, $\tilde{\mathbf{A}}_{t}=\hat{\mathbf{A}}-f(t,\delta)||\omega_{t}||_{\Sigma_{t}^{-1}}$
^$\hat{\mathcal{F}}_{t}$ Estimated feasible set w.r.t pessimistic estimate $\tilde{\mathbf{A}}$ at time $t$, =$\mathcal{F}=\{\tilde{\mathbf{A}}_{t}\in\mathbb{R}^{N\times K}:\tilde{\mathbf{A}}_{t}\boldsymbol{\pi}\leq 0\}$
$\mathcal{\mathbf{A}}$ The action set of $K$ possible choices
$\omega_{t}$ Allocation chosen at time $t$
@@
$N$ Number of Constraints
$\Gamma$ Slack of the optimisation problem
@@
$T$ Time Horizon
$r_{t},c_{t}$ Reward and cost observed at time $t$
@@
$l_{t}$ The Lagrangian multiplier at time $t$
@@
$\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ Set of alternative (confusing) instances for bandit instance ${\boldsymbol{\mu}}$
$\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})$ Estimated set of alternative (confusing) instances for bandit instance ${\boldsymbol{\mu}}$
@@
$\boldsymbol{\pi}^{\star}_{\mathcal{F}}$ True optimal policy w.r.t actual constraint set $\mathbf{A}$
@@
$\mathfrak{s}$ Shadow price $=\mathrel{\triangleq}\frac{\Gamma_{\mathrm{max}}}{\Gamma_{\mathrm{min}}}$
Appendix B Additional discussion on problem setting
@@ B.1 Extended related work
Existing work on adapting known constraints. In Multi-armed bandit literature, people often introduce constraints as a notion of safety where they impose known constraints on the chosen arm or on the exploration process. Wang et al., 2021b considers pure strategy (only one co-ordinate as chosen action) and imposes a safety threshold on the linear cost feedback of the chosen arm. On the other hand, the setting considered in Carlsson et al., (2024) is closer as it tracks an optimal policy w.r.t to a known set of known constraints. On the other hand, Liu et al., (2021) (Improvement over Pacchiano et al., (2020) in MAB setting) generalized the known constraint regret minimization setting by assuming existence of a set of general constraints. Our work captures the hardness of not knowing the constraint set while tracking the lower bound and also in sample complexity upper bounds of Algorithm 1 and 2. Our work also introduce shadow price as a novel term in pure exploration literature which characterises the extra cost that arises due to tracking the unknown constraints.
@@
Transductive Linear Bandit. In this setting Fiez et al., (2019) Camilleri et al., 2022b studies this setting with unknown linear constraints where we have to find the best safe arm in a finite set $\mathcal{Z}$ different than actual arm set $\mathcal{A}$. Our setting generalises the setting in the sense that the finite feasible set $\mathcal{Z}$ is not static, rather we track $\mathcal{Z}_{t}$ per time step $t\in\mathbb{N}$ and explore within that set to find the optimal allocation. At the end of exploration after hitting the stopping criterion at $\tau_{\delta}$ the agent recommends the optimal policy inside the set $\mathcal{Z}_{\tau_{\delta}}$.
@@
B.2 Motivations: Reductions to and generalisations of existing settings
@@
Thresholding Bandits. Our setting encompasses the thresholding bandit problem (Aziz et al., 2021a, ). Thresholding bandit is motivated from the safe dose finding problem in clinical trials, where one wants to identify the highest dose of a drug that is below a known safety level. This has also motivated the studies on safe arm identification (Wang et al., 2021b, ). Our setting generalises it further to detect the dose of the drug with highest efficacy while it is still below the safety level. We can formulate it as identifying $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}$, such that $\mathbf{I}\boldsymbol{\pi}\leq\mathbf{I}{\boldsymbol{\theta}}$. Rather, generalising the classical thresholding bandits, our formulation can further model the safe doses for the optimal cocktail of drugs, and $\theta$ can have different values across drugs, i.e we can consider different thresholds for different drugs.
@@
Optimal policy under Knapsack. Bandits under knapsack constraints have been studied both in best-arm identification (Li et al.,, 2023; Tran-Thanh et al.,, 2012; Li et al.,, 2021) and regret minimisation (Badanidiyuru et al.,, 2018; Agrawal and Devanur,, 2016; Immorlica et al.,, 2022; Agrawal and Devanur,, 2014; Agrawal et al.,, 2016; Sankararaman and Slivkins,, 2018; Ma,, 2014) literature. BAI under knapsacks is motivated by the fact that detecting an optimal arm might have additional resource constraints in addition to the number of required samples. This has led to study of BAI with knapsacks only under fixed-budget settings (Li et al.,, 2023). But as in regret-minimisation literature (Sankararaman and Slivkins,, 2018; Ma,, 2014), one might want to recommend a policy that maximises utility while satisfying knapsack constraints. For example, we want to manage caches where the recommended memory allocation should satisfy a certain resource budget. Thus, the recommended policy has to satisfy $\boldsymbol{\pi}^{\star}_{\tau}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in C_{\mathbf{A}}}\hat{{\boldsymbol{\mu}}}_{\tau_{\delta}}^{T}\boldsymbol{\pi}$, where $C_{\mathbf{A}}\mathrel{\triangleq}\{\mathbf{A}\boldsymbol{\pi}_{\tau_{\delta}}\leq c\}$. Naturally, this is a special case of our problem setting.
@@
Feasible arm selection. We look at the pure exploration problem of feasible arm selection studied by Katz-Samuels and Scott, (2018). Here, we think of a problem of workers having a multi-dimension vector representation where each index denotes the accuracy of that worker being able to identify a specific class label in a classification task in hand. The problem turns to be a feasible arm selection from a simple BAI problem when we impose a feasibility constraint that for example, the chosen worker should show more than 90% accuracy across all labels. We can generalise this setting in the sense that we are now not looking for a specific worker, rather we want to make a team of workers that has the highest utility. The recommended policy at time $t\in\mathbb{N}$, $\max_{\boldsymbol{\pi}\in\Delta_{K-1}}{\boldsymbol{\mu}}_{t}^{T}\boldsymbol{\pi}$ such that $f^{T}\boldsymbol{\pi}\geq\tau$ where $\tau$ is the desired threshold level. The generalisation of the setting pitch in as thresholds of $\tau$ can have different values corresponding to different workers.
@@
BAI with fairness across sub-populations. The Best Arm Identification with fairness Constraints on Sub-population (BAICS) studied in Wu et al., (2023) aims on selecting an arm that must be fair across all sub-populations rather than the whole population in standard BAI setting. Let, there are $\boldsymbol{l}$ sub-populations and $\mu_{a}$ are the means corresponding to the a-th arm. Finding only the optimal arm $K_{\mathrm{BAI}}=\operatorname*{arg\,max}_{k\in[{K}]}{\boldsymbol{\mu}}_{k}$ may not be enough because it may not perform equally good for all the $l$ sub-populations. Then the arm should belong to a set $C:=\{k\in[{K}]|{\boldsymbol{\mu}}_{k,m}\geq 0,m\in[l]\}$ where the observation for arm k and population m comes from $\mathcal{N}(\mu_{k,m},1)$ It ensures that the chosen arm does not perform too bad for any sub-population. Let us think of a problem where there are $l$ sub-groups of patients and we have ${K}$ number of drugs to administer with reward means $\mu_{k},k\in[{K}]$. We are looking for a combination of drugs rather than a single drug to administer as $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\pi\in\Delta_{K}}{\boldsymbol{\mu}}^{T}\pi$ such that $\mathds{1}_{{\boldsymbol{\mu}}_{m}\geq 0}^{T}\boldsymbol{\pi}=1,\forall m\in[l]$. Thus, BAICS is a special case of ours.

### v2

@@ 4 Algorithm Design
11:  Update loss for regret minimizer: Update with $L_{t}=\langle\boldsymbol{\omega}_{t}^{\star},U_{t}\rangle-{\boldsymbol{l}_{t}^{\star}}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}^{\star}$
12:end while
13:Recommended policy: $\hat{{\boldsymbol{\pi}}}=\operatorname*{arg\,max}\limits_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\hat{{\boldsymbol{\mu}}}_{t}^{\top}\boldsymbol{\pi}$
@@
LAGEX. Algorithms based on track and stop mechanism tend to fail in case of larger problems where efficient optimization becomes a challenge due to the use of a max-min oracle per step. To improve on this, we leverage the two-player zero sum game approach introduced in Degenne et al., 2019b . Algorithm 3 also starts by playing each arm once to warm up parameter estimates. Then it uses a allocation player (We have used AdaGrad) to optimize the allocation $\boldsymbol{\omega}_{t}$ in Line 5 against the most confusing instance w.r.t current estimate of $\boldsymbol{\mu}$} optimised by a instance player which minimizes $\sum_{a=1}^{K}\boldsymbol{\omega}_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-{\boldsymbol{l}_{t}^{\star}}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$ with respect to ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}({\boldsymbol{\mu}},\boldsymbol{\pi})$. Since our search space in closed and convex, the allocation player enjoys sub-linear regret of $\mathcal{O}(\sqrt{t\log t})$, whereas the instance player computes the best confusing instance using Proposition 1. Then in Line 11, Adagrad loss function is updated with a loss by introducing optimism as $U_{t}$ defined in Line 10. Rest of the mechanism goes as usual
Theorem 7 (Asymptotic Optimality of LAGEX).
Under Assumption 1 and 2, the expected sample complexity of LAGEX satisfies $\lim_{\begin{subarray}{c}\delta\rightarrow 0\end{subarray}}\frac{\mathbb{E}[\tau_{\delta}]}{\log(1/\delta)}\leq T_{\mathcal{F},r}({\boldsymbol{\mu}}).$
Detailed proofs of this section are in Appendix G.
@@ 5 Experimental Analysis
Now, we empirically test performance of proposed algorithms and the baselines. We refer to Appendix I for additional experiments and other details necessary for reproducibility. Code is available at this Link. [^2] Baselines: “CTnS-WLag” and “CGE-WLag” is the CTnS and CGE algorithm of Carlsson et al., (2024) under unknown constraints without Lagrangian relaxation. “PTnS” is Projected Track and Stop algorithm.
Synthetic Data: Setup 1 We evaluate with two environments having means $1.5,1.0,\mu_{3},0.4,0.3,0.2,0.1]$.
Observation 1: Universality. We vary $\mu_{3}$ from $0.5$ to $2.5$. For each environment, we plot the corresponding unconstrained BAI lower bounds (in red) and lower bounds under constraints (in blue) in Figure 5. We observe that the constraint problem gets easier with increasing $\mu_{3}$. In contrast, the BAI problem changes non-monotonically. BAI problem gets harder when $\mu_{3}$ is around $1.5$ as the suboptimality gap gets very small. But the constraint problem stays easier than BAI. In Figure 5, we also plot the median sample complexity of LAGEX across these environments over 500 runs. LAGEX grows parallel to the lower bound under constraints and can track it across environments.
Observation 2: Efficiency. We run all algorithms in two environment: (i) hard with $\mu_{3}=0.5$ and (ii) easy with $\mu_{3}=1.3$. We call the first environment hard as it is harder than solving BAI and similarly, the second environment easy. In Figure 5 and 7 we observe that (i) among the algorithms with unknown constraints LAGEX incur the least sample complexity, and (ii) we pay a minimal cost than the known constraint Lagrangian algorithms in hard env whereas the price of estimating constraints is prominent in easy env.
Figure 4: Lower bounds with and without constraints, and LAGEX for $\mu_{3}\in[0.5,2.5]$ in Setup 1.
[…]
Real Data: IMDB-50K Dataset We evaluate LATS, LAGEX, and the baselines on IMDB 50K dataset (Maas et al.,, 2011). For ease of comparison, we use the same bandit environment as in (Carlsson et al.,, 2024). We use 12 movies. We search for the optimal policy which allocates weight at most $0.3$ to action movies and at least $0.3$ to family and drama movies.
@@
6 Discussion & Future Work
We study pure exploration under unknown linear constraints. We encode the coupled effect of estimating both mean vector and constraints via a Lagrangian relaxation of the lower bound for known constraints. We further design an optimistic estimate of the feasible set to ensure identification of the $r$-optimal feasible policy- leading to two algorithms, LATS and LAGEX. We prove their sample complexity upper bounds and conduct numerical experiments to observe that LAGEX is the most efficient among baselines.
From the lower bound perspective, one might be interested to derive the lower bounds when both constraints and rewards are considered as feedback. Algorithmically, it would be intriguing to extend our Lagrangian technique to non-linear constraints.
Acknowledgments
We acknowledge the ANR JCJC project REPUBLIC (ANR-22-CE23-0003-01), the PEPR project FOUNDRY (ANR23-PEIA-0003), and the Inria-ISI Kolkata associate team SeRAI for supporting the project. We also thank Achraf Azize for interesting discussions.
@@ Appendix A Notations
K{K} Number of Arms
AA True constraint set
@@
$\mathcal{F}$ True feasible set w.r.t $A$, $\mathcal{F}=\{A\in\mathbb{R}^{d\times K}:A\boldsymbol{\pi}\leq 0\}$
[…]
$\hat{\mathcal{F}}_{t}$ Estimated feasible set w.r.t pessimistic estimate $\tilde{A}$ at time $t$, $\mathcal{F}=\{\tilde{A}_{t}\in\mathbb{R}^{d\times K}:\tilde{A}_{t}\boldsymbol{\pi}\leq 0\}$
$\mathcal{A}$ The action set of $K$ possible choices
${\boldsymbol{\omega}}_{t}$ Allocation chosen at time $t$
@@
NN Number of Constraints
$\Gamma$ Slack of the optimisation problem, i.e., $\Gamma\mathrel{\triangleq}\min_{i\in[d]}(-A\boldsymbol{\pi}^{\star}_{\mathcal{F}})$
@@
$r_{t},\mathbf{c}_{t}$ Reward and cost observed at time $t$
@@
$\boldsymbol{l}_{t}$ The Lagrangian multiplier at time $t$
@@
$\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})$ Set of alternative (confusing) instances for bandit instance $\boldsymbol{\mu}$}
$\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})$ Estimated set of alternative (confusing) instances for bandit instance $\boldsymbol{\mu}$}
@@
$\boldsymbol{\pi}^{\star}_{\mathcal{F}}$ True optimal policy w.r.t actual constraint set AA
@@
$\mathfrak{s}$ Shadow price $\mathrel{\triangleq}\frac{\Gamma_{\mathrm{max}}}{\Gamma_{\mathrm{min}}}$
Table 1: Summary of Notations
Appendix B Discussion on The Problem Setting and Connection to Literature
@@ B.1 Extended Related Work
Existing work on adapting known constraints. In Multi-armed bandit literature, people often introduce constraints as a notion of safety where they impose known constraints on the chosen arm or on the exploration process. Wang et al., 2021b considers pure strategy (only one co-ordinate as chosen action) and imposes a safety threshold on the linear cost feedback of the chosen arm. On the other hand, the setting considered in Carlsson et al., (2024) is closer as it tracks an optimal policy w.r.t to a known set of known constraints. On the other hand, Liu et al., (2021) (Improvement over Pacchiano et al., (2020) in MAB setting) generalized the known constraint regret minimization setting by assuming existence of a set of general constraints. Our work captures the hardness of not knowing the constraint set while tracking the lower bound and also in sample complexity upper bounds of Algorithm 2 and 3. Our work also introduce shadow price as a novel term in pure exploration literature which characterises the extra cost that arises due to tracking the unknown constraints.
@@
Transductive Linear Bandit. In this setting formalised by Fiez et al., (2019), Camilleri et al., 2022b studies this setting with unknown linear constraints where we have to find the best safe arm in a finite set $\mathcal{Z}$ different than actual arm set $\mathcal{A}$. Our setting generalises the setting in the sense that the finite feasible set $\mathcal{Z}$ is not static, rather we track $\mathcal{Z}_{t}$ per time step $t\in\mathbb{N}$ and explore within that set to find the optimal allocation. At the end of exploration after hitting the stopping criterion at $\tau_{\delta}$ the agent recommends the optimal policy inside the set $\mathcal{Z}_{\tau_{\delta}}$.
@@
B.2 Motivations: Reductions to and Generalisations of The Existing BAI Settings
@@
Thresholding Bandits. Our setting encompasses the thresholding bandit problem (Aziz et al., 2021a, ). Thresholding bandit is motivated from the safe dose finding problem in clinical trials, where one wants to identify the highest dose of a drug that is below a known safety level. This has also motivated the studies on safe arm identification (Wang et al., 2021b, ). Our setting generalises it further to detect the dose of the drug with highest efficacy while it is still below the safety level. We can formulate it as identifying $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$, such that $I\boldsymbol{\pi}\leq I{\boldsymbol{\theta}}$. Rather, generalising the classical thresholding bandits, our formulation can further model the safe doses for the optimal cocktail of drugs, and $\theta$ can have different values across drugs, i.e we can consider different thresholds for different drugs.
@@
Optimal policy under Knapsack. Bandits under knapsack constraints have been studied both in best-arm identification (Li et al.,, 2023; Tran-Thanh et al.,, 2012; Li et al.,, 2021) and regret minimisation (Badanidiyuru et al.,, 2018; Agrawal and Devanur,, 2016; Immorlica et al.,, 2022; Agrawal and Devanur,, 2014; Agrawal et al.,, 2016; Sankararaman and Slivkins,, 2018; Ma,, 2014) literature. BAI under knapsacks is motivated by the fact that detecting an optimal arm might have additional resource constraints in addition to the number of required samples. This has led to study of BAI with knapsacks only under fixed-budget settings (Li et al.,, 2023). But as in regret-minimisation literature (Sankararaman and Slivkins,, 2018; Ma,, 2014), one might want to recommend a policy that maximises utility while satisfying knapsack constraints. For example, we want to manage caches where the recommended memory allocation should satisfy a certain resource budget. Thus, the recommended policy has to satisfy $\boldsymbol{\pi}^{\star}_{\tau}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in C_{A}}\hat{{\boldsymbol{\mu}}}_{\tau_{\delta}}^{\top}\boldsymbol{\pi}$, where $C_{A}\mathrel{\triangleq}\{A\boldsymbol{\pi}_{\tau_{\delta}}\leq c\}$. Naturally, this is a special case of our problem setting.
@@
Feasible arm selection. We look at the pure exploration problem of feasible arm selection studied by Katz-Samuels and Scott, (2018). Here, we think of a problem of workers having a multi-dimension vector representation where each index denotes the accuracy of that worker being able to identify a specific class label in a classification task in hand. The problem turns to be a feasible arm selection from a simple BAI problem when we impose a feasibility constraint that for example, the chosen worker should show more than 90% accuracy across all labels. We can generalise this setting in the sense that we are now not looking for a specific worker, rather we want to make a team of workers that has the highest utility. The recommended policy at time $t\in\mathbb{N}$, $\max_{\boldsymbol{\pi}\in\Delta_{K-1}}{\boldsymbol{\mu}}_{t}^{\top}\boldsymbol{\pi}$ such that $f^{\top}\boldsymbol{\pi}\geq\tau$ where $\tau$ is the desired threshold level. The generalisation of the setting pitch in as thresholds of $\tau$ can have different values corresponding to different workers.
@@
BAI with fairness across sub-populations. The Best Arm Identification with fairness Constraints on Sub-population (BAICS) studied in Wu et al., (2023) aims on selecting an arm that must be fair across all sub-populations rather than the whole population in standard BAI setting. Let, there are $\boldsymbol{l}$ sub-populations and $\mu_{a}$ are the means corresponding to the a-th arm. Finding only the optimal arm $K_{\mathrm{BAI}}=\operatorname*{arg\,max}_{k\in[{K}]}{\boldsymbol{\mu}}_{k}$ may not be enough because it may not perform equally good for all the $l$ sub-populations. Then the arm should belong to a set $C:=\{k\in[{K}]|{\boldsymbol{\mu}}_{k,m}\geq 0,m\in[l]\}$ where the observation for arm k and population m comes from $\mathcal{N}(\mu_{k,m},1)$ It ensures that the chosen arm does not perform too bad for any sub-population. Let us think of a problem where there are $l$ sub-groups of patients and we have K{K} number of drugs to administer with reward means $\mu_{k},k\in[{K}]$. We are looking for a combination of drugs rather than a single drug to administer as $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\pi\in\Delta_{K}}{\boldsymbol{\mu}}^{\top}\pi$ such that $\mathds{1}_{{\boldsymbol{\mu}}_{m}\geq 0}^{\top}\boldsymbol{\pi}=1,\forall m\in[l]$. Thus, BAICS is a special case of ours.